if you randomly select 2 cards from a standard 52-card deck. find the probability that the first card and the second card is red if
a: you replace the first card before selecting the second
b: you do not replace the first card

Answer by stanbon(48568)

if you randomly select 2 cards from a standard 52-card deck. find the probability that the first card and the second card is red if
a: you replace the first card before selecting the second
Ans: If you replace the card the events are independent; that means what
happened on the 1st draw has no effect on what happens on the 2nd draw.
In that case P(red and red)= p(red)*P(red) = 1/2*1/2 = 1/4=0.25
-----------------
b: you do not replace the first card
Now the events are dependent because you are drawing the 2nd card from
a "different" deck; the deck is not missing one red card.
P(red and red) = p(red)*P(red|red) = (26/52) * (25/51) = (1/2)(25/51)
=25/102=0.245098...
----------
Cheers,
Stan H.